A wave with a wavelength of 1.5 m and a speed of 60.0 m/s Find the frequency of the wave.


1
Expert's answer
2021-04-07T10:09:10-0400
"v=f\lambda,""f=\dfrac{v}{\lambda}=\dfrac{60\ \dfrac{m}{s}}{1.5\ m}=40\ Hz."